OpenStudy (anonymous):

The diameter of a basketball rim is 18 inches. A standard basketball has a circumference of 30 inches. About how much room is there between the ball and the rim in a shot in which the ball goes in exactly in the center of the rim?

OpenStudy (anonymous):

4.225

OpenStudy (anonymous):

8.45inches

OpenStudy (anonymous):

For the ball: C = 2 x pi x r, so r = C/(2 x pi) = 30 / (2 x 3.14) = 4.8 in. diameter of ball = 2 x r = 9.6 in.